OpenStudy (anonymous):

Solve the system by using a matrix equation. 2x - 3y = 3 5x - 7y = 9

OpenStudy (phi):

can you set up the matrix ?

OpenStudy (phi):

it would be 2 x 2 matrix

OpenStudy (anonymous):

-5 (2x-3y=3) 2 (5x-7y=9) -10x+15y=-15 10x-14y=18

OpenStudy (anonymous):

Now we have : y=3

OpenStudy (phi):

once you set up the matrix, call it A you have the problem A v = b where v is the vector [ x y] and b= [3 9] transposed (vertical) use http://www.mathsisfun.com/algebra/matrix-inverse.html and scroll down to 2 x 2 matrix to find how to invert A A_inv * A v = A_inv b v = A_inv b

OpenStudy (anonymous):

x = 3 Understand ?! :D

OpenStudy (phi):

E.ali, the question is asking to use a matrix approach, presumably to learn how to use matrices

OpenStudy (anonymous):

I'm not sure how I would set up the matrix from the equations